how do I make this exercise -5n-8(1+7n)=-8
-8(-8x-6)=-6x-22

Answer by mananth(16946)

how do I make this exercise

-5n-8(1+7n)=-8
-5n-8-56n=-8
-61n=8-8
-61n=0
n=0
...
-8(-8x-6)=-6x-22
64x+48=-6x-22
add 6x to both sides
64x+6x+48=-22
70x+48=-22
add -48 to both sides
70x=-22-48
70x=-70
/70
x=-1
...
m.ananth@hotmail.ca